OpenStudy (anonymous):

solve the equation using the quadratic formula. then solve the equation by factoring to check your solution(s). 3x^2+7x-24=13x please explain

OpenStudy (turingtest):

which part do you have problems with, the factoring or the quadratic formula?

OpenStudy (lilg132):

do you know the quadratic formula?

OpenStudy (anonymous):

yes. I need help on the factoring part

OpenStudy (lilg132):

ok well just to make sure you do know you have to make the equation = 0

OpenStudy (anonymous):

so, 3x^2-6x-24=0

OpenStudy (lilg132):

yes

OpenStudy (anonymous):

now how do you factor with a number other than 1 for a

OpenStudy (lilg132):

a would always be the number e.g. 3x in one bracket and x in the other to give it 3x^2

OpenStudy (anonymous):

(3x- )(x+ ) and the blanks have to add to 6? or 24?

OpenStudy (lilg132):

yes have to add to -6x and multiply to be + 24 (remember - multiplied by - gives +)

OpenStudy (lilg132):

i've learnt that the only way to factorise is through trial and error unless someone else knows a different way

OpenStudy (anonymous):

so I need to find factors of 24 that subtract to 6…right?

OpenStudy (lilg132):

yes you need to find factors of 24 that add to give you -6

OpenStudy (anonymous):

would (3x-12)(x+6) work?

OpenStudy (anonymous):

because you have to multiply by the 3 too

OpenStudy (lilg132):

- 12 * 6 would not give you 24

OpenStudy (lilg132):

sorry -24 is what you need

OpenStudy (lilg132):

i misread the equation

OpenStudy (anonymous):

but 3x times x=3x^2, 3x times 6=18x, -12 times x=-12x, and -12 times 6 =…oh nevermind

OpenStudy (lilg132):

haha

OpenStudy (lilg132):

but i can say your on the right tracks try and change the number you have in the second bracket and it should fix it

OpenStudy (lilg132):

sorry - 12 * something should give - 24

OpenStudy (anonymous):

-2

OpenStudy (lilg132):

+2

OpenStudy (anonymous):

2

OpenStudy (lilg132):

yes

OpenStudy (lilg132):

try that

OpenStudy (anonymous):

(3x-12)(x+2)=3x^2+6x-12x-24=3x^2-6-24

OpenStudy (lilg132):

-6x not -6 but yes thats correct

OpenStudy (anonymous):

thanks so much. i have a quiz on this tomorrow and i didn't get it, but now i do! Oh right, 6x.

OpenStudy (lilg132):

your welcome :p dont forget to click good answer ;)